# RAD: Towards Trustworthy Retrieval-Augmented Multi-modal Clinical Diagnosis

### v1

@@ 1 Introduction
The rapid development of multimodal learning [45, 36] has revolutionized numerous fields by enabling models to process and integrate diverse data types, including images, texts, audio, and structured records [4, 76, 11]. Biomedical applications particularly benefit from these advancements, given that diagnostic workflows inherently depend on multimodal evidence, ranging from radiographic imaging and reports to electronic health records (EHR) [51, 74, 16]. For instance, radiologists integrate X-ray or MRI scans with textual pathology reports, while clinicians combine electronic health records, vital signs, and even genomic data to form comprehensive patient profiles. Accordingly, recent research efforts have increasingly focused on developing multimodal architectures tailored to healthcare challenges, seeking to enhance diagnostic precision through cross-modal synergy [67, 31, 60, 6]. While these approaches demonstrate significant progress in integrating data from different modalities, they often overlook the foundational principles governing clinical decision-making.
@@
Medical analysis fundamentally differs from natural scene understanding in its strict adherence to evidence-based principles, relying heavily on structured protocols [30, 49]. Clinical decisions must be grounded in standardized diagnostic criteria derived from patient-specific symptoms, imaging findings, and laboratory results. This inherent rigor poses a critical challenge for black-box neural networks, whose vague decision-making mechanisms hinder trustworthy and practical deployment in clinical settings [52, 48]. Consequently, there has been growing interest in integrating medical knowledge into AI models to simultaneously improve model performance and interpretability [8, 72, 9].
@@
Existing approaches primarily focus on knowledge injection during pretraining phases. Researchers enhance the text encoders by pretraining them on large-scale medical corpora [29, 46] or leveraging structured knowledge graphs to imbue models with semantic relationships between biomedical entities [47, 35]. While effective in expanding the semantic coverage of text encoders, these approaches often struggle to explicitly integrate fine-grained knowledge tailored for downstream diagnostic tasks. To this end, we argue that effective knowledge integration requires task-centric, holistic alignment with disease-level knowledge throughout the entire diagnostic pipeline. As illustrated in Figure 1(a), our framework systematically integrates refined knowledge to guide input augmentation, feature extraction, and modality fusion, contrasting with prior methods confined to a single perspective. Figure 1(b) presents a case of the model’s attention distribution over the input text. The previous model fails to concentrate on critical indicators, but focuses on obvious disease terms in the reports. While the RAD model can not only attend to these terms but also consider other guideline-recommended key indicators. The explicit knowledge guidance enables RAD to prioritize critical indicators tailored for the current disease, making trustworthy diagnoses aligned with clinical standards.
@@
  • We aligned MIMIC-CXR [27] and MIMIC-IV [28] to construct the MIMIC-ICD53 dataset, covering three modalities with 53 types of disease. Extensive experiments on our dataset and three other public datasets demonstrate the superiority of RAD over SOTA baselines across various metrics.
@@ 2.1 Multimodal Learning in Medicine
Recent years have witnessed significant advancements in the field of multimodal learning, with models such as CLIP [45], BLIP [32], and LLaVA [36] exhibiting remarkable capabilities in natural domains. These developments have spurred increasing interest in extending multimodal frameworks to the medical field, where the integration of diverse data modalities demonstrates prominent potential in diagnostic tasks. Current research focus lies in multimodal pretraining methods, which focus on cross-modal alignment between imaging and textual data to improve the representation transferability [7, 17]. ConVIRT [71] and GLoRIA [25] pioneered the application of CLIP-style architectures in the medical domain by constructing image-text pairs from radiology datasets. MedCLIP [55] and BiomedCLIP [69] addressed the scarcity of paired medical image-text data by leveraging multi-source datasets, achieving state-of-the-art performance. Beyond pretraining methods, multimodal fusion approaches aim at integrating information from different modalities for diagnostic applications [19, 56, 65]. MedFuse [20] introduced an LSTM-based temporal fusion method of time-series data and X-ray images. HEALNet [22] proposed a hybrid early-fusion method to learn from data sources with different structures. While these works have made significant strides, they often operate without explicit guidance from medical knowledge when addressing specific diagnosis tasks. In contrast, considering the evidence-based nature of medicine [50, 43], RAD explicitly incorporates task-specific knowledge to guide both representation extraction and multimodal fusion processes.
@@ 2.2 Medical Knowledge Injection
Injecting professional knowledge into AI models is a prevalent strategy to improve their domain-specific capabilities [39, 64]. Various techniques have been investigated to incorporate medical knowledge into the models. Pretraining-based approaches train the text encoder on extensive medical domain corpora, such as PubMedBERT [18] and HUATUO-GPT [68]. Other methods like KAD [70] and DRAGON [66] leverage structured knowledge graphs of medical entities for pre-training to enhance the text encoder’s comprehension of medical terminology. While showing empirical effectiveness, these knowledge integration methods remain primarily confined to the pre-training phase, providing only implicit guidance during the subsequent diagnostic stage. With the rapid development of large language models (LLMs) [1, 53, 34], various Retrieval-Augmented Generation (RAG) methods have been proposed to enhance the generation process of medical LLMs [13, 60]. These methods dynamically retrieve external medical knowledge to improve the performance of LLMs on question-answering (QA) tasks [59, 33]. Building upon this foundation, multimodal RAG approaches further retrieve similar data samples (e.g., image-report pairs) for visual question-answering (VQA) [73, 58]. In contrast to RAG methods that online retrieve knowledge to augment input for generative QA/VQA tasks, our framework adopts a structured approach for discriminative tasks by performing offline retrieval of disease-specific knowledge, which is systematically incorporated to guide model training.
@@ 3.2.1 Guideline Retrieval and Refinement
For a given dataset with $m$ diseases, our objective is to retrieve the most relevant knowledge from the knowledge corpus $P$, including but not limited to: associated symptoms, imaging characteristics, and critical examination/laboratory indicators. We adopt MedCPT [26], a dual-encoder model optimized for medical scenarios, as the retriever. Specifically, the article encoder $R_{A}(\cdot)$ is utilized to convert the corpus $P$ into dense vectors for retrieval. The disease names $E$ are used as the input query of the query encoder $R_{Q}(\cdot)$. The obtained embeddings are then used to calculate the similarity as $\operatorname{Sim}(E,P)={R}_{Q}(E)^{\top}{R}_{A}(P)$. For each disease with a name $e_{i}\in E$, we preserve the top-$k$ retrieved documents as:
@@
where $c_{i,j}\in\mathcal{C}_{i}$ is the $j$-th document. This process yields standardized, well-structured diagnostic guidelines that preserve critical clinical information while eliminating noise and redundancy. In practice, we choose Qwen2.5-72B [63] as the LLM. Examples of the guideline and prompt templates are provided in Appendix B.2.
@@ 4 Experiments
4.1 Experimental Setup
Table 1: Detailed information of the datasets.
Dataset Anatomy Modality Label Sample
MIMIC-ICD53 Chest X-ray Image & Report & EHR (Lab Results) 53 51830
Harvard-FairVLMed Eye Fundus Image & Report & Demographics 1 10000
SkinCAP Skin Dermatology Image & Report 50 2526
NACC Brain 3D MRI Image & EHR (Lab Results) 11 4199
Datasets. We evaluate RAD on four multimodal medical datasets with different anatomies, including MIMIC-ICD53, Harvard-FairVLMed [40], SkinCAP [75], and NACC [5]. MIMIC-ICD53 is constructed through the alignment and integration of MIMIC-CXR [27] and MIMIC-IV [28], comprising chest X-ray images, corresponding reports, and EHRs, annotated with 53 diseases under the ICD [42] standard. For laboratory indicators in the EHR, we quantified the numerical results on a scale of 1 to 10 based on the upper and lower limits of their normal range. We will release the dataset on PhysioNet [41]. Details of dataset construction are provided in Appendix C.1.1. Harvard-FairVLMed, SkinCAP, and NACC are multimodal datasets focusing on eyes, skin, and brain, respectively. All patient data has been de-identified. More detailed statistics of datasets are presented in Table 1.
Baselines. We select representative baseline methods in the medical field, including large-scale pre-training model BiomedCLIP [69], knowledge-enhanced pre-training method KAD [70], and state-of-the-art multimodal fusion methods MedFuse [20], DrFuse [65], and HEALNet [22].
Implementation Details. In practice, Top-$k$ in Eq.(1) is set to 10. All guidelines obtained by Eq.(2) and the indicators used in Algorithm 1 are manually verified to avoid potential factual errors. The sample ratio of negative samples in $\mathcal{L}_{\text{GECL}}$ is set to 5. The default backbone of the text encoder and image encoder is ClinicalBERT [54] and ResNet-50 [21], respectively. The hyperparameters $\alpha$ and $\beta$, which serve as the balancing ratio between different losses, are set to be 1e21e-2 and 1e11e-1, respectively. All experiments are conducted on a single NVIDIA A100 GPU.
@@ 4.3.1 Interpretability from Visual Perspective
To quantitatively assess the impact of knowledge injection from the visual perspective, we conduct zero-shot grounding experiments on the ChestX-Det dataset [37]. The results shown in Table 3 demonstrate a significant improvement in mIoU scores for lesion detection after the injection of refined guidelines. Besides, Figure 3 illustrates multiple cases of lesion grounding. For clearer visualization, we overlay spectrum heatmaps on the original CXR images, together with the ground truth bounding box highlighted in red. A comparison between the lesions identified by the model and the bounding boxes marked by clinical experts reveals a notable improvement in alignment when our guidelines are applied. This indicates that the model’s focus is more accurately directed toward clinically significant lesions, emphasizing RAD’s enhanced diagnosis capabilities and interpretability under the guidance of external knowledge.
4.3.2 Interpretability from Textual Perspective
Symmetrically, we calculate the guideline recall defined in Section 3.3.1 to investigate the effect of our guideline injection from the textual perspective. As presented in Table 4, incorporating knowledge via RAD prominently increases the recall value from 24.76% to 65.62%. This indicates that RAD indeed injects guideline-derived knowledge into the model, thereby enhancing its focus on key information mentioned in the guideline. Notably, the conventional model exhibits extremely low recall (<10%) on Alanine Aminotransferase (ALT) and Aspartate Aminotransferase (AST). This may stem from their inability to understand highly specialized, rare medical terms. In contrast, RAD explicitly highlights the importance of these indicators in the guideline, leading to significant recall improvement. This finding underscores the necessity of flexible knowledge adaptation for downstream tasks rather than static pretraining paradigms. Overall, the enhanced guideline recall demonstrates that RAD enables the model to make reliable evidence-based diagnoses according to guidelines. This improvement also aligns with the qualitative attention patterns observed in Figure 1. To further substantiate the interpretability of RAD, we provide more and clearer visualization cases in Appendix C.5.
@@ 4.4 Ablation Study
To demonstrate the robustness and flexibility of our method, we verify RAD on different encoder backbone combinations. Specifically, we iteratively replaced the default visual encoder and text encoder with two other popular architectures, ViT [15] and BioClinicalBERT [2]. As illustrated in Figure 4, RAD consistently offers substantial performance gain across all four combinations of backbones. This not only highlights the insensitivity of our approach to different backbone architectures but also underscores its robustness and generalizability. Specifically, ResNet and ViT exhibit comparable performance gains, while ClinicalBERT shows more pronounced improvement than BioClinicalBERT. Overall, RAD exhibits the robust ability to generalize to diverse data and model structures, ensuring reliable performance in various scenarios. Detailed results, more ablation studies, and hyperparameter analysis are presented in Appendix C.6.
Figure 5: Performance comparison with MLLMs. We convert the single-label dataset FairVLMed into the visual question answering format and evaluate MLLMs under both few-shot and supervised fine-tuning (SFT) settings.
@@ 4.5 Discussion
Multimodal large language models exhibit remarkable capabilities in visual content understanding and generalization. To further validate the effectiveness of RAD, we compare with state-of-the-art MLLMs, including Qwen2.5-VL-7B [3], HuatuoGPT-Vision-7B [10], and Lingshu-7B [61]. As presented in Figure 5, our discriminative framework achieves superior performance with significantly lower computational cost. These results demonstrate that complex diagnostic tasks are better suited for specialized discriminative models than generative MLLMs. The significant performance gap, observed on the simplest single-label dataset, underscores the practical advantages of our approach in clinical applications where both accuracy and efficiency are critical.

### v2

@@ 1 Introduction
The rapid development of multimodal learning [radford2021learning, liu2023visual] has revolutionized numerous fields by enabling models to process and integrate diverse data types, including images, texts, audio, and structured records [baltruvsaitis2018multimodal, zhou2024reprogramming, chen2023vast]. Biomedical applications particularly benefit from these advancements, given that diagnostic workflows inherently depend on multimodal evidence, ranging from radiographic imaging and reports to electronic health records (EHR) [suk2014hierarchical, zhou2022generalized, dou2020unpaired]. For instance, radiologists integrate X-ray or MRI scans with textual pathology reports, while clinicians combine electronic health records, vital signs, and even genomic data to form comprehensive patient profiles. Accordingly, recent research efforts have increasingly focused on developing multimodal architectures tailored to healthcare challenges, seeking to enhance diagnostic precision through cross-modal synergy [yunflex, li2024lorkd, xu2024ram, bernal2017deep]. While these approaches demonstrate significant progress in integrating data from different modalities, they often overlook the foundational principles governing clinical decision-making.
@@
Medical analysis fundamentally differs from natural scene understanding in its strict adherence to evidence-based principles, relying heavily on structured protocols [lekadir2021future, stiglic2020interpretability]. Clinical decisions must be grounded in standardized diagnostic criteria derived from patient-specific symptoms, imaging findings, and laboratory results. This inherent rigor poses a critical challenge for black-box neural networks, whose vague decision-making mechanisms hinder trustworthy and practical deployment in clinical settings [sun2020saunet, salahuddin2022transparency, han2025trustworthy]. Consequently, there has been growing interest in integrating medical knowledge into AI models to simultaneously improve model performance and interpretability [chen2021learning, zhang2017mdnet, chen2024cod].
@@
Existing approaches primarily focus on knowledge injection during pretraining phases. Researchers enhance the text encoders by pretraining them on large-scale medical corpora [lee2020biobert, rasmy2021med] or leveraging structured knowledge graphs to imbue models with semantic relationships between biomedical entities [rotmensch2017learning, liu2021auto]. While effective in expanding the semantic coverage of text encoders, these approaches often struggle to explicitly integrate fine-grained knowledge tailored for downstream diagnostic tasks. To this end, we argue that effective knowledge integration requires task-centric, holistic alignment with disease-level knowledge throughout the entire diagnostic pipeline. As illustrated in Figure 1(a), our framework systematically integrates refined knowledge to guide input augmentation, feature extraction, and modality fusion, contrasting with prior methods confined to a single perspective. Figure 1(b) presents a case of the model’s attention distribution over the input text. The previous model fails to concentrate on critical indicators, but focuses on obvious disease terms in the reports. While the RAD model can not only attend to these terms but also consider other guideline-recommended key indicators. The explicit knowledge guidance enables RAD to prioritize critical indicators tailored for the current disease, making trustworthy diagnoses aligned with clinical standards.
@@
  • We aligned MIMIC-CXR [johnson2019mimic] and MIMIC-IV [johnson2023mimic] to construct the MIMIC-ICD53 dataset, covering three modalities with 53 types of disease. Extensive experiments on our dataset and three other public datasets demonstrate the superiority of RAD over SOTA baselines across various metrics.
@@ 2.1 Multimodal Learning in Medicine
Recent years have witnessed significant advancements in the field of multimodal learning, with models such as CLIP [radford2021learning], BLIP [li2023blip], and LLaVA [liu2023visual] exhibiting remarkable capabilities in natural domains. These developments have spurred increasing interest in extending multimodal frameworks to the medical field, where the integration of diverse data modalities demonstrates prominent potential in diagnostic tasks. Current research focus lies in multimodal pretraining methods, which focus on cross-modal alignment between imaging and textual data to improve the representation transferability [boecking2022making, eslami2023pubmedclip]. ConVIRT [zhang2022contrastive] and GLoRIA [huang2021gloria] pioneered the application of CLIP-style architectures in the medical domain by constructing image-text pairs from radiology datasets. MedCLIP [wang2022medclip] and BiomedCLIP [zhang2023biomedclip] addressed the scarcity of paired medical image-text data by leveraging multi-source datasets, achieving state-of-the-art performance. Beyond pretraining methods, multimodal fusion approaches aim at integrating information from different modalities for diagnostic applications [harutyunyan2019multitask, wolf2022daft, yao2024drfuse]. MedFuse [hayat2022medfuse] introduced an LSTM-based temporal fusion method of time-series data and X-ray images. HEALNet [hemker2024healnet] proposed a hybrid early-fusion method to learn from data sources with different structures. While these works have made significant strides, they often operate without explicit guidance from medical knowledge when addressing specific diagnosis tasks. In contrast, considering the evidence-based nature of medicine [subbiah2023next, peng2023ai], RAD explicitly incorporates task-specific knowledge to guide both representation extraction and multimodal fusion processes.
@@ 2.2 Medical Knowledge Injection
Injecting professional knowledge into AI models is a prevalent strategy to improve their domain-specific capabilities [liu2021finbert, yang2024textbook]. Various techniques have been investigated to incorporate medical knowledge into the models. Pretraining-based approaches train the text encoder on extensive medical domain corpora, such as PubMedBERT [gu2021domain] and HUATUO-GPT [zhang2023huatuogpt]. Other methods like KAD [zhang2023knowledge] and DRAGON [yasunaga2022deep] leverage structured knowledge graphs of medical entities for pre-training to enhance the text encoder’s comprehension of medical terminology. While showing empirical effectiveness, these knowledge integration methods remain primarily confined to the pre-training phase, providing only implicit guidance during the subsequent diagnostic stage. With the rapid development of large language models (LLMs) [achiam2023gpt, touvron2023llama, liu2024deepseek], various Retrieval-Augmented Generation (RAG) methods have been proposed to enhance the generation process of medical LLMs [chen2025towards, xu2024ram]. These methods dynamically retrieve external medical knowledge to improve the performance of LLMs on question-answering (QA) tasks [xiong2024benchmarking, li2025biomedrag]. Building upon this foundation, multimodal RAG approaches further retrieve similar data samples (e.g., image-report pairs) for visual question-answering (VQA) [zhao2023retrieving, xia2024mmed]. In contrast to RAG methods that online retrieve knowledge to augment input for generative QA/VQA tasks, our framework adopts a structured approach for discriminative tasks by performing offline retrieval of disease-specific knowledge, which is systematically incorporated to guide model training.
@@ 3.2.1 Guideline Retrieval and Refinement
For a given dataset with $m$ diseases, our objective is to retrieve the most relevant knowledge from the knowledge corpus $P$, including but not limited to: associated symptoms, imaging characteristics, and critical examination/laboratory indicators. We adopt MedCPT [jin2023medcpt], a dual-encoder model optimized for medical scenarios, as the retriever. Specifically, the article encoder $R_{A}(\cdot)$ is utilized to convert the corpus $P$ into dense vectors for retrieval. The disease names $E$ are used as the input query of the query encoder $R_{Q}(\cdot)$. The obtained embeddings are then used to calculate the similarity as $\operatorname{Sim}(E,P)={R}_{Q}(E)^{\top}{R}_{A}(P)$. For each disease with a name $e_{i}\in E$, we preserve the top-$k$ retrieved documents as:
@@
where $c_{i,j}\in\mathcal{C}_{i}$ is the $j$-th document. This process yields standardized, well-structured diagnostic guidelines that preserve critical clinical information while eliminating noise and redundancy. In practice, we choose Qwen2.5-72B [yang2024qwen2] as the LLM. Examples of the guideline and prompt templates are provided in Appendix B.2.
@@ 4 Experiments
4.1 Experimental Setup
Table 2: Detailed information of the datasets.
Dataset Anatomy Modality Label Sample
MIMIC-ICD53 Chest X-ray Image & Report & EHR (Lab Results) 53 51830
Harvard-FairVLMed Eye Fundus Image & Report & Demographics 1 10000
SkinCAP Skin Dermatology Image & Report 50 2526
NACC Brain 3D MRI Image & EHR (Lab Results) 11 4199
Datasets. We evaluate RAD on four multimodal medical datasets with different anatomies, including MIMIC-ICD53, Harvard-FairVLMed [luo2024fairclip], SkinCAP [zhou2024skincap], and NACC [beekly2007national]. MIMIC-ICD53 is constructed through the alignment and integration of MIMIC-CXR [johnson2019mimic] and MIMIC-IV [johnson2023mimic], comprising chest X-ray images, corresponding reports, and EHRs, annotated with 53 diseases under the ICD [world2009international] standard. For laboratory indicators in the EHR, we quantified the numerical results on a scale of 1 to 10 based on the upper and lower limits of their normal range. We will release the dataset on PhysioNet [moody2022physionet]. Details of dataset construction are provided in Appendix C.1.1. Harvard-FairVLMed, SkinCAP, and NACC are multimodal datasets focusing on eyes, skin, and brain, respectively. All patient data has been de-identified. More detailed statistics of datasets are presented in Table 2.
Baselines. We select representative baseline methods in the medical field, including large-scale pre-training model BiomedCLIP [zhang2023biomedclip], knowledge-enhanced pre-training method KAD [zhang2023knowledge], and state-of-the-art multimodal fusion methods MedFuse [hayat2022medfuse], DrFuse [yao2024drfuse], and HEALNet [hemker2024healnet].
Implementation Details. In practice, Top-$k$ in Eq.(1) is set to 10. All guidelines obtained by Eq.(2) and the indicators used in Algorithm 1 are manually verified to avoid potential factual errors. The default backbone of the text image encoder is ClinicalBERT [wang2023optimized] and ResNet-50 [he2016deep], respectively. The hyperparameters $\alpha$ and $\beta$, which serve as the balancing ratio between different losses, are set to be 1e21e-2 and 1e11e-1, respectively. All experiments are conducted on a single NVIDIA A100 GPU.
@@ 4.3 Interpretability Evaluation
4.3.1 Interpretability from Textual Perspective
To quantitatively assess the impact of knowledge injection from the textual perspective, we calculate the guideline recall defined in Section 3.3.1. As presented in Table 4, incorporating knowledge via RAD prominently increases the recall value from 24.76% to 65.62%. This indicates that RAD indeed injects guideline-derived knowledge into the model, thereby enhancing its focus on key information mentioned in the guideline. Notably, the conventional model exhibits extremely low recall (<10%) on Alanine Aminotransferase (ALT) and Aspartate Aminotransferase (AST). This may stem from their inability to understand highly specialized, rare medical terms. In contrast, RAD explicitly highlights the importance of these indicators in the guideline, leading to significant recall improvement. This finding underscores the necessity of flexible knowledge adaptation for downstream tasks rather than static pretraining paradigms. Overall, the enhanced guideline recall demonstrates that RAD enables the model to make reliable evidence-based diagnoses according to guidelines. This improvement also aligns with the qualitative attention patterns observed in Figure 1. To further substantiate the interpretability of RAD, we provide more and clearer visualization cases in Appendix C.5.
@@ 4.3.2 Interpretability from Visual Perspective
Symmetrically, we conduct zero-shot grounding experiments on the ChestX-Det dataset [liu2020chestx]. The results shown in Table 3 demonstrate a significant improvement in mIoU scores for lesion detection after the injection of refined guidelines. Besides, Figure 3 illustrates multiple cases of lesion grounding. For clearer visualization, we overlay spectrum heatmaps on the original CXR images, together with the ground truth bounding box highlighted in red. A comparison between the lesions identified by the model and the bounding boxes marked by clinical experts reveals a notable improvement in alignment when our guidelines are applied. This indicates that the model’s focus is more accurately directed toward clinically significant lesions, emphasizing RAD’s enhanced diagnosis capabilities and interpretability under the guidance of external knowledge.
@@ 4.4 Ablation Study
To demonstrate the robustness and flexibility of our method, we verify RAD on different encoder backbone combinations. Specifically, we iteratively replaced the default visual encoder and text encoder with two other popular architectures, ViT [dosovitskiy2020image] and BioClinicalBERT [alsentzer2019publicly]. As illustrated in Figure 4, RAD consistently offers substantial performance gain across all four combinations of backbones. This not only highlights the insensitivity of our approach to different backbone architectures but also underscores its robustness and generalizability. Specifically, ResNet and ViT exhibit comparable performance gains, while ClinicalBERT shows more pronounced improvement than BioClinicalBERT. Overall, RAD exhibits the robust ability to generalize to diverse data and model structures, ensuring reliable performance in various scenarios. Detailed results, more ablation studies, and hyperparameter analysis are presented in Appendix C.6.
Figure 5: Performance comparison with MLLMs. We convert the single-label dataset FairVLMed into the VQA format and evaluate MLLMs under both few-shot and supervised fine-tuning settings.
@@ 4.5 Discussion
Multimodal large language models exhibit remarkable capabilities in visual content understanding and generalization. To further validate the effectiveness of RAD, we compare with state-of-the-art MLLMs, including Qwen2.5-VL-7B [bai2025qwen2], HuatuoGPT-Vision-7B [chen2024huatuogpt], and Lingshu-7B [xu2025lingshu]. As presented in Figure 5, our discriminative framework achieves superior performance with significantly lower computational cost. These results demonstrate that complex diagnostic tasks are better suited for specialized discriminative models than generative MLLMs. The significant performance gap, observed on the simplest single-label dataset, underscores the practical advantages of our approach in clinical applications where both accuracy and efficiency are critical.